OpenStudy (anonymous):

Jackie has two jobs. She earns $14/h in Job A and $16/h in Job B. She would like to earn $560 in a week. Let a represent the number of hours she works in Job A, and let b represent the number of hours she works in Job B. Which function relates the number of hours she works in Job B to the number of the hours she works in Job A?

OpenStudy (anonymous):

@satellite73 @fallenangelorchid

OpenStudy (fallenangelorchid):

let 'a' be the number of hours @ $14/hr 'b' be the number of hours @ $16/hr Since she needs $560 then 14a + 16b = 560 ---------------------- To get the value of b in terms of a, transpose 14a to the right side 16b = -14a + 560 Divide the whole equation with 16 b = -14/16 a + 560/16 ---------------------- You reduce that and you will have the answer

OpenStudy (anonymous):

how do you reduce that

OpenStudy (fallenangelorchid):

To make this easier its b=8/7a - 35

OpenStudy (anonymous):

how did you get 35

OpenStudy (anonymous):

and wouldn't it be 7/8

OpenStudy (anonymous):

@fallenangelorchid